OpenStudy (anonymous):

Combine the like terms to simplify the expression: 4 + 7xy2 + 5y3 + 10x –4xy2 – 8x – 4y3 + 9 A. y6 + 3x2y2 + 10x + 13 B. y3 + 3xy2 + 2x + 13 C. y3 + 3xy2 + 10x + 13 D. y3 – 3xy2 + 2x + 9

OpenStudy (anonymous):

@mathway @200205650

OpenStudy (anonymous):

constants: 4+9 = 13 xy^2: 7-4 = 3xy^2 y^3: 5-4 = y^3 x: 10-8 = 2x y3 + 3xy2 +2x +13 (answer B) am i right

OpenStudy (anonymous):

Look at all the terms to see which terms happen more than once. The + or - sign in front of the term tells you whether it's positive or negative. Then add together all the sets of terms which are the same. Put the term with the highest exponent first, then put in the rest of the exponents in descending order. 4+9= 13 7xy2+(-4xy2)= 3xy2 5y3+(-4y3)= y3 10x+(-8x)= 2x 13+3xy2+y3+2x y3+3xy2+2x+13

OpenStudy (anonymous):

it's c

OpenStudy (anonymous):

c y3+3xy2+2x+13

OpenStudy (anonymous):

oh do you guys agree?

OpenStudy (purplemexican):

i get 3xy2+y3+2x+13

OpenStudy (anonymous):

You're correct. It's B

OpenStudy (anonymous):

yeah c y3 + 3xy2 +2x +13

OpenStudy (anonymous):

yea B

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

OH LOL

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

yeah i got y3 + 3xy2 +2x +13 and saw c right away xD

OpenStudy (purplemexican):

i got what looks like c in an odd order 3xy^2+y^3+2x+13

OpenStudy (anonymous):

lol